GYROIDS OF CONSTANT MEAN CURVATURE

KARSTEN GROSSE-BRAUCKMANN

Abstract. We use Brakke’s Surface Evolver to deform a triply periodic minimal surface, the gyroid, into a continuous family of constant mean curvature surfaces with the same symmetry. We discuss stability and bifurcation problems for these surfaces.